# Domain Adaptation-Enhanced Searchlight: Enabling brain decoding from visual perception to mental imagery

### v1

@@ Appendix A Supplementary Results
Figure 10: Subject variability of the within-domain searchlight accuracy. For each subject, the color of each voxel represents the average perception-to-perception accuracy over the 100 repetitions of the experiment in the regions where such accuracy is statistically significantly above 0.5.
@@
Figure 11: Subject variability of the cross-domain searchlight accuracy. For each subject, the color of each voxel represents the average baseline perception-to-imagery accuracy over the 100 repetitions of the experiment in the regions where such accuracy is statistically significantly above 0.5.
@@
Figure 12: Subject variability of the cross-domain searchlight accuracy with 100 target domain instances available during training. For each subject, the color of each voxel represents the average accuracy over the 100 repetitions of the experiment in the regions where such accuracy is statistically significantly above 0.5.
Appendix B Brief description of each DA method used
  • 1.
@@ Appendix B Brief description of each DA method used
    BalancedWeighting [20]: A supervised, instance-based method that fits the base estimator $h$ source and target labeled data according to the modified loss $(1-\gamma)\mathcal{L}(h(X_{s}),y_{s})+\gamma\mathcal{L}(h(X_{t}),y_{t})$, where the parameter $\gamma$ controls the ratio between fitting the source data and adapting to the target data, with higher values indicating more adaptation. In our case, $\mathcal{L}$ is the binary cross-entropy and $\gamma=0.5$. BalancedWeighting can be used with any supervised estimator.
  • 2.
@@
  • 8.
    NearestNeighborsWeighting [69]: An unsupervised, instance-based approach that computes, for each source domain sample, the number of neighbours inside a certain radius in the target data set. Source instances are reweighted proportionately.
  • 9.

### v3

@@ Appendix A Supplementary Results
Figure 11: Subject variability of the within-domain searchlight accuracy. For each subject, the colour of each voxel represents the average perception-to-perception accuracy over the 100 repetitions of the experiment in the regions where such accuracy is statistically significantly above 0.5.
@@
Figure 12: Subject variability of the cross-domain searchlight accuracy. For each subject, the colour of each voxel represents the average baseline perception-to-imagery accuracy over the 100 repetitions of the experiment in the regions where such accuracy is statistically significantly above 0.5.
@@
Figure 13: Subject variability of the cross-domain searchlight accuracy with 100 target domain instances available during training. For each subject, the colour of each voxel represents the average accuracy over the 100 repetitions of the experiment in the regions where such accuracy is statistically significantly above 0.5.
Appendix B Results on “Generic Object Decoding” dataset
B.1 Dataset description
The publicly available “Generic Object Decoding” dataset [35] gathers fMRI scans from 5 healthy subjects that were exposed to visual stimuli from 50 different categories (perception condition) and were later instructed to imagine those categories (imagery condition). Both of these experiments constituted the evaluation data for the methodology proposed in the original paper [35]. Data acquisition and preprocessing are discussed in [35].
In the perception condition, each category was presented a total of 35 times and the experiment involved 35 runs. The stimuli presentation order was randomized across runs. Each category included a single representative image, which was extracted from the THINGS [59] database.
The imagery condition involved 20 runs, each of which included half of the categories in a semi-random order, such that every two consecutive runs contained all the categories. Each run was composed of 25 imagery events, and therefore, each category was imagined a total of 10 times during the experiment.
This dataset allows the opportunity to test the applicability of DA methods to assess the transferability of visual perception decoding to the imagery domain in a multiclass classification setting. However, the reduced amount of imagery data per class (10 examples) is a key limitation for training and evaluation, both of which require all categories to be present in the corresponding data partitions. Furthermore, the need to ensure that examples from the same experimental run are not present in both train and test partitions renders training impossible given the distribution of categories across the runs in the data acquisition step.
For the reasons listed above, investigating the impact of DA on this dataset is only feasible if the number of classes is reduced in a way that preserves the correlation between class labels and neural activity. We propose a semantic grouping into 9 broader classes (Table 4), one of which must be discarded due to insufficient imagery data (“Weapons”, containing only the subclass “Cannon”: 10 examples). This semantic grouping results in an imbalanced dataset, because each broader class contains a different number of subclasses. Many other semantic groupings are possible, as long as they are crafted in such a way that preserves the human interpretation of the stimuli.
Broad Category Original Subclasses Number of Examples
Animals Goldfish, Owl, Common iguana, Duck, Swan, Conch, Crab, Killer whale, Leopard, Bat, Housefly, Butterfly, Goat, Camel, Domestic llama 150
Clothing and Accessories Cowboy hat, Silk hat, Football helmet, Sock, Welder’s mask 50
Household Items Beer mug, Bowling ball, Hammock, Mailbox, Umbrella 50
Musical Instruments Electric guitar, Grand piano, Harp, Mandolin, Tambourine 50
Structures and Monuments Coffin, Gravestone, Minaret, Stained-glass window 40
Technology and Devices IPod, Microwave, Videocassette recorder, Washing machine 40
Tools and Equipment Fire extinguisher, Knob, Tweezer, Planchet, Shredder 50
Vehicles and Transportation Airliner, Barrow, Bulldozer, Canoe, Covered wagon, Snowmobile 60
Weapons Cannon 10
Table 4: Broad categories, their original subclasses, and the number of examples.
B.2 Machine learning task and validation framework
We performed the comparison of the 15 DA methods under study, as well as the NAIVE and BASELINE methods as defined in Section 4. The validation framework was exactly the same, except for the number of data partitions, which decreased from 100 to 24 as a consequence of requiring all 8 classes to be present for both training and testing. The number of target domain instances in the training set ($N_{t}$) was also increased to account for the multiclass setting ($N_{t}=200,250,300$). Random over-sampling was used in both the source and target domains to balance the dataset before training.
We used all the voxels from the largest ROI provided by the dataset authors, which covers the Visual Cortex (VC) as defined in [35] Supplementary Figure 1. Because of the reduced number of subjects and viable data partitions, the total experiments per algorithm were significantly fewer for the “Generic Object Decoding” dataset, impacting the power of the statistical tests.
B.3 Comparison of DA techniques
From Figure 14 we conclude that PRED, FA, BW and TAB were the best algorithms. Among those, PRED was statistically significantly better than NAIVE ($\alpha=0.05$).
Figure 15 indicates that PRED significantly outperformed other algorithms 47 times, and was only outperformed once (by NNW). It was the most outperforming algorithm. Similarly, FA, BW and TAB were outperformed only once (by NNW).
Due to the decrease in the number of subjects and feasible data partitions, the total number of experiments per algorithm was dramatically lower for the “Generic Object Decoding” dataset, which affects the power of the statistical tests summarized in figures 14 and 15. Figure 16 shows the results for each subject, revealing that FA and PRED were better than NAIVE in 4 subjects, BW for 3 and TAB for 2.
Figure 15: Frequency table. For each cell $(i,j)$, the numerical value shows for how many subjects algorithm $i$ was significantly superior to algorithm $j$. The sum of each row accounts for the number of times each algorithm was significantly better than others, and the sum of each column shows how many times each algorithm was significantly worse than others.
Figure 16: Critical Difference diagrams for each subject, comparing different DA techniques in the Visual Cortex.
Appendix C DA-enhanced searchlight results with the Balanced Weighting technique
Figure 17: Brain regions where BW decoding is significantly better than the naïve approach ($p<0.05$, whole-brain corrected).
As illustrated in Supplementary Figure 18, the cross-subject average accuracy achieved using BW significantly exceeded chance levels across a broader range of brain regions compared to RTLC. Notably, with BW, our DA-enhanced searchlight successfully classifies imagery of Living versus Non-Living items in regions that extend into the frontal lobe. However, in terms of symmetry between the left and right hemispheres, the accuracy map for BW exhibits less balance compared to RTLC. This asymmetry is likely attributable to cross-subject averaging, as indicated by the comparison between Figure 9 and Supplementary Figure 19. Specifically, BW demonstrates poor decoding performance for subjects 9, 15, and 17, whereas RTLC achieves successful decoding for these individuals.
Figure 18: Between-subject average of the balanced accuracy of RTLC and BW in the regions that obtained statistically above chance decoding with each corresponding DA technique.
Figure 19: Inter-subject variability of the DA-enhanced searchlight accuracy. For each subject, the colour of each voxel represents the average BW accuracy over the 100 repetitions of the experiment in the regions where such accuracy is statistically above 0.5.
Appendix D Brief description of each DA method used
@@
  • BalancedWeighting [20]: A supervised, instance-based method that fits the base estimator $h$ source and target labelled data according to the modified loss $(1-\gamma)\mathcal{L}(h(X_{s}),y_{s})+\gamma\mathcal{L}(h(X_{t}),y_{t})$, where the parameter $\gamma$ controls the ratio between fitting the source data and adapting to the target data, with higher values indicating more adaptation. In our case, $\mathcal{L}$ is the binary cross-entropy and $\gamma=0.5$. BalancedWeighting can be used with any supervised estimator.
@@
  • NearestNeighborsWeighting (NNW) [76]: An unsupervised, instance-based approach that computes, for each source domain sample, the number of neighbours inside a certain radius in the target data set. Source instances are reweighted proportionately.
@@
  • FineTuning (FT) [27]: A feature-based, supervised DA method originally developed for Convolutional Neural Networks, that reuses the internal layers of a network trained in the source domain and furthers training of the external layers using a limited amount of target domain data.
  • DeepCORAL (Deep CORrelation ALignment - DCORAL) [31]: Correlation Alignment is a feature-based and unsupervised DA method that aligns the second-order statistics of the source and target distributions using a linear transformation. Similarly, Deep CORAL is a nonlinear extension that aligns correlations of layer activations in deep neural networks.
  • DANN (Discriminative Adversarial Neural Network) [30]: DANN is a feature-based DA method, which strives to find a new representation of the input features in which source and target data are indistinguishable by any discriminator network. This new representation is given by an encoder which is learned along the discriminator using a reversal gradient layer. In parallel, a task network is learned on the encoded space.
  • MCD (Maximum Classifier Discrepancy) [29]: MCD is a feature-based, unsupervised DA method that looks for a new representation of the input features which minimizes the discrepancy between the source and target domains. The discrepancy is computed through adversarial training of three networks: An encoder and two classifiers, each of which learn the task on the source and the target domains. A reversal layer is placed between the encoder and the two classifiers to perform adversarial training.